OpenStudy (anonymous):

How do I take efficient fast math note in computer? Is there any software or any strategies that will help to take perfect math note fast or learn thing and keep thing remember

OpenStudy (anonymous):

Microsoft Math is a free one.

OpenStudy (e.mccormick):

For perfect look to the notes there is \(\LaTeX\), which has many free implementations. Even online ones. For just the math part of this, http://www.codecogs.com/latex/eqneditor.php may work. For memorization, there are many methods of mnemonics that can help. Be sure to back them up with self quizzing.

OpenStudy (anonymous):

Thanks so much for help!!! :)

OpenStudy (anonymous):

But I am looking software

OpenStudy (e.mccormick):

\(\LaTeX\) is also what is used to make math show up here. \(\int\limits_x^a42\text{da}\)

OpenStudy (e.mccormick):

I have a posting that includes the software. http://openstudy.com/study#/updates/52bd214ae4b01cdca44a70d2

OpenStudy (e.mccormick):

Here is an example of some stuff I used it for. This is part of a document I was doing on Linear Algebra.

OpenStudy (e.mccormick):

To start, I would try the online version to see if you like it or not: https://www.writelatex.com/ http://www.scribtex.com/ https://www.sharelatex.com/ If it works for you, then I would install something like MiKTEX so you can just do it offline, etc.

OpenStudy (anonymous):

miktex is a free ?

OpenStudy (anonymous):

ok cool

OpenStudy (e.mccormick):

Yep!

OpenStudy (anonymous):

Thanks :) I am struggling with math note

OpenStudy (anonymous):

I cant remember or organize them. I needed that will help me to make perfect math note fast and learn the concept better

OpenStudy (e.mccormick):

Yah, I have dysgraphia so my handwriting is really bad. So writing it is not always great. With this I can type in \(\dfrac{-b\pm\sqrt{b^2-4ac}}{4a}\) and it comes out clear.

OpenStudy (anonymous):

Wow! great :)

OpenStudy (e.mccormick):

I did that for my linear class. Took my scribbles from class and turned them into a document.

OpenStudy (anonymous):

so I have to start with online sites that you gave then later i can install miktex

OpenStudy (e.mccormick):

You can go right to miktex if you want. However, by trying it online you can see if it works or you without installing anything.

OpenStudy (anonymous):

So in sharetex can I write equation as well?

OpenStudy (e.mccormick):

Yes. And the codecogs editor can let you visually edit things. Sometimes I use it for a document when I want to proof it visually before I put it in.

OpenStudy (anonymous):

ok, writelatex is not free.

OpenStudy (anonymous):

ok thanks alot for help! :) I try scribtex or sharetex

OpenStudy (e.mccormick):

Ah, thought their personal one was free. I see I am wrong.

OpenStudy (anonymous):

thats fine, I will try sharetex or scribtex

OpenStudy (e.mccormick):

Yah, and it looks like those are merging. They used to be seperate. Hehe.

OpenStudy (anonymous):

Ok, so with this I can take quickly good math note and save it. Right?

OpenStudy (e.mccormick):

Quick depends on how good you are at typing. They will be good. And you can save them.

OpenStudy (anonymous):

so ok, I am not quite fast in typing. But I do want to do something that help me to write perfect math note from my long chapters faster and quick

OpenStudy (e.mccormick):

Well, what I typed above was: `\dfrac{-b\pm\sqrt{b^2-4ac}}{4a}` and that made \(\dfrac{-b\pm\sqrt{b^2-4ac}}{4a}\).

OpenStudy (anonymous):

:( really struggling with it because, in all those long math chapter eating up my times in note and end up disorganized and lost

OpenStudy (anonymous):

It happen with me most of the time, which disappointed me so much

OpenStudy (anonymous):

Anyways, thanks a lot for help!!! :)

OpenStudy (e.mccormick):

No problem. I hope it works well for you!

OpenStudy (anonymous):

Me too, I will let you know if it works great! :) I hope it will!

OpenStudy (e.mccormick):

Kk. Have fun!

OpenStudy (anonymous):

Thanks again for awesome help!!! :)